# DeepVigor+: Scalable and Accurate Semi-Analytical Fault Resilience Analysis for Deep Neural Networks

### v1

@@ III-A Fault Model
Transistor miniaturization leads to more efficient computational digital systems, whereas its reliability side-effects are becoming increasingly profound. With transistor scaling soft error rates increase significantly posing reliability concerns, in particular in the deployment of safety-critical applications [41, 42, 43]. Single Event Upset (SEU) due to soft errors may affect memory cells and flip a bit in a stored value. SEU in DNN accelerators affects either the parameters of a DNN (i.e., weights) or its activations (i.e., layers’ inputs), which are stored in memory elements such as registers or on-chip memories [16, 22].
For reliability analysis, the multi-bit fault model is more realistic, however, it requires a huge fault space to consider all combinations, i.e., $3^{n-1}$ combinations where n is the number of bits. On the other hand, it is demonstrated that the single-bit fault model provides a high fault coverage of multiple-bit faults [44]. Therefore, analyzing the reliability of DNNs based on a single-bit fault model is valid for obtaining the VF of models. This fault model is in line with the other works in the literature [23, 25].
Therefore, in this paper we build our analysis based on the single-bit fault model in weights and activations separately. We assume that a single parameter or a single input to a layer is faulty which is propagated to the output of the corresponding layer. By mathematical analysis, we model the fault propagation to the output of neurons and DNNs, resulting in calculating VF for them.
III-B Single Fault Analysis in 32-bit Floating-Point
To model the behavior of faults for the analysis, we assume that at an inference, a single fault may influence the value of a single input activation to a neuron or a single parameter in a CNN. First, we analyze the effect of a single bitflip on the 32-bit floating-point data representation to comprehend how a value might change due to a single fault. IEEE-754 32-bit floating-point data type is shown in Fig. 3. It contains 1 sign bit, 8 exponent bits, and 23 fraction bits, and represents a number based on Eq. (2).
$number=(-1)^{sign}\times 2^{E-127}\times(1+\sum_{i=0}^{i=23}{b_{23-i}\times 2^{i}})$ (2)
@@ III-B Single Fault Analysis in 32-bit Floating-Point
In this regard, we consider the following lemmas:
  • Lemma 1: Any bitflip from 1 to 0 in a value decreases the value as $\epsilon$ while $\epsilon<0$, and any bitflip from 0 to 1 increases it as $\epsilon$ while $\epsilon>0$.
@@
      Sign: approximate $\epsilon$ as ±$\pm 2^{log_{2}(2\times x)}$.
@@
      Exponent: approximate $\epsilon$ as ±$\pm 2^{i-23}$, assuming that $x$ is a small value.
@@
      Mantissa: approximate $\epsilon$ as ±$\pm 2^{i}$.
    In all cases, $\epsilon$ can be approximated as the nearest value of the power of two to the actual $\epsilon$. This approximation can lead to a unified representation for $\epsilon$. To that end, $\epsilon$ might be negative or positive (based on Lemma 1) and might be small or big (based on Lemma 2). Eventually, we can express it as a unified representation as shown in Eq. (4).
@@
  • Lemma 4: When a faulty value is used in a multiplication operation in CNNs, the erroneous output can also be approximated. It has been observed that the parameters in CNNs are mainly distributed around 0 and in the range of $[-1,1]$ [47]. In order to approximate the error of multiplying two values when one of them is faulty, we can analyze it based on Eq. (5), where $x$ and $y$ are fault-free values and $x^{\prime}$ is the faulty value after a bitflip in $x$.
@@
III-C Single Fault Error Propagation in DNNs
We analyze the single bitflip error propagation in CNNs considering their effect on the values of numbers. Each neuron in a convolutional (CONV) layer operates as shown in Eq. (6), where the Output Feature Map (OFMap) in $l$th layer and $k$th channel is obtained by the summation of multiplications between weights ($\hat{w}$) and Input Feature Map (IFMap, $\hat{x}$) plus bias ($b$). In CONV layers, $\hat{w}$ and $\hat{x}$ are 3-Dimensional (3D) $c_{in}\times n\times n$ arrays, where $c_{in}$ is the number of inputs channels to the layer, and $n$ is the kernel size.
@@ III-C Single Fault Error Propagation in DNNs
Figure 4: Fault propagation in a CNN in the case of a single bitflip in a weight.
@@
Based on the aforementioned theoretical analysis, to analyze the effect of single faults on CNNs, regardless of the fault occurrence in activations or weights, we only need to identify an added value $\delta$ (as a power of two) at the output of the target neuron/channel where it misclassifies the golden class of the CNN for each input image.
Figure 5: An overview of the conducted steps in DeepVigor+.
@@ III-D DeepVigor+: Fault Resilience Analysis for DNNs
In this section, the steps of DeepVigor+ are presented. Fig. 5 illustrates the steps of the DeepVigor+ methodology for fault resilience analysis of DNNs against single faults. The method’s objective is to provide the Vulnerability Factor (VF) for layers and the entire DNN model when a single fault happens in activations or weights.
The inputs of the method are a pre-trained CNN and validation data in a dataset. During the analysis, an OFMap or output channel is targeted, and multiple steps are taken to provide the VF metrics:
@@
    Vulnerability Values Identification: obtains vulnerability values for the target OFMap/channel
@@
By obtaining the VF for the analyzed neurons and filters in a DNN, the VF for layers and the entire DNN can be derived. The details of each step are explained in the following. Noteworthy that each step presents a detailed description for analyzing a target neuron and then briefly links it to weights in filters.
Step 1 - Single Bitflip Analysis: As mentioned above, single faults at the inputs of a neuron are considered. This step conducts bitflips in the input activations of the target neuron to construct Vulnerability Values Search Space (VVSS) and Error Distribution Map (EDM) for the target neuron. VVSS is a set of candidate values representing all possible $\delta$ in Eq. (9). In other words, VVSS represents output errors produced by single bitflips in any bit locations of the inputs of the target neuron. EDM represents the distribution of $\delta$ in Eq. (9) throughout the approximated $\delta$ values with a power of two.
Obtaining VVSS and EDM requires performing a single bitflip for each input and multiplying it by its corresponding weight as shown in Eq. (8). Nonetheless, it is an exhaustive operation with high time complexity. This complexity can be remarkably reduced by leveraging Algorithm 1. In this algorithm, for each bitflip, we obtain all possible $\delta$ produced at the output of the neuron for all inputs at once.
Since each input might be faulty separately, to produce all errors in a neural operation, we first flip the $i$th bit in all inputs (line 3) and then convert the binary representation to a value (line 4). Then, the difference of the erroneous_inputs with the fault-free input is computed (line 5), which represents all possible errors at inputs ($\epsilon$ in Eq. (8)) for all inputs added by a single bitflip in either of them. Then, a point-wise multiplication between the erroneous_inputs and weights results in producing all possible errors ($\delta$ in Eq. (9)) at the output (line 6).
Algorithm 1 Error Analysis for a Neuron
@@
Algorithm 1 produces all errors at the output of the target neuron. Thereafter, based on the presented theory in the subsections III-B and III-C, we generate the distribution of all produced errors as an Error Distribution Map (EDM) based on the Candidate Vulnerability Values (CVVs) which are a set of numbers as a power of two, shown in Eq. (11). Based on the experimental observations, we limit the CVV in the analysis between $-2^{10}$ and $2^{10}$ for large values. The error values between $-2^{-10}$ and $2^{-10}$ are merged as their propagation effects on CNNs are negligible and masked.
@@
EDM contains the distribution ratio of the existing values over the values in output_errors_list in Algorithm 1, between each consecutive values in CVV, as shown in Eq. (12). EDM provides the ratio of error distribution with respect to each candidate vulnerability value, demonstrating how much each CVV represents the produced errors at the output of the target neuron. Based on the EDM, each CVV whose distribution ratio is not zero, is added to the Vulnerability Value Search Space (VVSS). This means that VVSS contains a subset of CVV that are the approximated errors produced by single faults in the inputs and might lead to misclassification at the output of the CNN.
@@
Step 2: Vulnerability Values Identification: This step exploits VVSS to identify the vulnerability values for the target neuron (i.e., OFMap) by exploring its constructed VVSS. As mentioned, VVSS represents output errors induced by a single fault occurring at the inputs of the target neuron. The objective is to identify the maximum negative and minimum positive vulnerability values among the existing ones in VVSS for a neuron that misclassify the DNN’s outputs from its golden classification for each input data.
@@
    $VVSS_{[-\infty,-1]}$: contains CVVs between $[-\infty,1]$ that have a non-zero distribution ratio.
@@
    $VVSS_{(-1,0)}$: contains CVVs between $(-1,0)$ that have a non-zero distribution ratio.
@@
    $VVSS_{[1,+\infty]}$: contains CVVs between $[1,+\infty]$ that have a non-zero distribution ratio.
@@
In box 3, a loss function $\mathcal{L}$ is calculated to obtain the effect of added $\delta$ to the target OFMap on the output classes based on the summation of the differences between all output class’s logits from the golden class. In this loss function, $\mathcal{E}_{c_{t}}$ represents the erroneous output logit of the golden top class and $\mathcal{E}_{c_{j}}$ is the erroneous output logit of any other class.
In box 4, the gradient of loss function $\mathcal{L}$ w.r.t. the target OFMap is calculated to check if the CNN might be misclassified or not, when the target OFMap is erroneous. If the gradient is 0, faults producing positive deviation in the corresponding neuron do not lead to misclassification. In this case, the neuron’s vulnerability value is considered the biggest CVV assumed value for positive numbers (i.e., $2^{10}$) in box 5, and the algorithm ends. If the gradient is not equal to 0, a value can be found for $\delta$ in the target OFMap which misclassifies the DNN. The rest of the algorithm attempts to find the minimum positive $\delta$ in VVSS.
In box 6, it is examined if the CNN misclassifies the input from its golden classification when $\delta=1$ for the target neuron, determining the initialization for $\delta$ in the next steps. If the input is misclassified when $\delta=1$, it means that its vulnerability value is less than $1$ and we should explore $VVSS_{(0,1)}$ (box 7), otherwise, $VVSS_{[1,+\infty)}$ should be explored (box 8). In the case of exploring $VVSS_{(0,1)}$, $\delta$ is set to the middle element of the set, and if it does not misclassify the golden output, the next bigger CVV should be explored. This process continues until the first value which misclassifies the output is found (loop 9). The final positive vulnerability value ($\delta^{+}$) for the target neuron is the minimum value in positive VVSS that does not misclassify the input for CNN from its golden classification.
To identify the maximum negative vulnerability value ($\delta^{-}$) that does not misclassify the input from its golden classification, the same procedure is conducted to explore negative values in VVSS. As a result, the Vulnerability Value Range (VVR) for the corresponding neuron is obtained and is expressed in Eq. (13). VVR represents all induced error values to the outputs of a neuron leading to a misclassification. Noteworthy that since bitflips in 32-bit floating-point might lead to large values, we assumed that any value larger than $2*{10}$ and less than $2^{-10}$ are critical for CNNs, therefore, these values represent +$\infty$+ ∞ and $\infty$- ∞ respectively.
@@
The step of vulnerability values identification for a target filter is similar to the one for a neuron. It is conducted similarly to the illustrated flowchart in Fig. 6 to obtain VVR for the corresponding output channel resulting in VVR for each target channel in a DNN.
Step 3: Obtaining Vulnerability Factors: As mentioned, an Error Distribution Map (EDM) is obtained for the target neuron in step 1, representing the distribution ratio for each vulnerability value. On the other hand, the Vulnerability Value Range (VVR) is obtained in step 2. This step aims to map VVR to EDM to provide the VF for the target neuron, expressing the probability of misclassification in the case of a single bitflip in its inputs.
Fig. 7 depicts how VF can be obtained by mapping VVR to aggregated EDM. Aggregated EDM represents the summation of the distribution ratio for all values less than a CVV. As shown in Fig. 7, all values between $(\delta^{-},\delta^{+})$ are non-vulnerable meaning that if an error deviates the OFMap as big as any value in this range, it will not misclassify the golden output. Otherwise, the error value is vulnerable leading to a misclassification.
VF is obtained by subtracting the distribution ratio of the minimum positive vulnerability value ($\delta{+}$) and the maximum negative vulnerability value ($\delta^{-}$) as shown in Eq. (14). This equation expresses what portion of all errors produced at the output are critical for the CNN in terms of misclassification which is equivalent to the probability that a fault misclassifies the CNN’s output.
@@
The VF for a filter is obtained similarly. the EDM and VVR for the target filter are obtained in the previous steps. In this step, they are exploited to provide the VF for the target filter. Deriving the VF for individual neurons and filters leads to obtaining the VF for CONV layers and the entire CNN model. Since the analysis is performed for neurons and filters separately, it can provide separate VF for layers and for the entire model based on filters and neurons. The Layer Vulnerability Factor (LVF) is the average of VF for all activations/filters, and the Model Vulnerability Factor (MVF) is the average of LVFs within it.
Furthermore, to obtain the total MVF of the entire model using the obtained detailed VFs based on the activations and filters analysis, Eq. (15) is introduced, where L is the total number of layers in a CNN, $N_{l}$ and $W_{l}$ are the total number of output activations and weights in layer $l$. This equation is a layerwise weighted average of LVFs throughout the layers of a CNN including both activations and filters vulnerability analysis, leading to the MVF for the entire model, representing the probability of misclassification if a fault occurs either in activations or weights.
@@
$\frac{\sum_{l=1}^{L}(\frac{N_{l}}{N_{l}+W_{l}}\times LVF_{act_{l}}+\frac{W_{l}}{N_{l}+W_{l}}\times LVF_{weight_{l}})}{L}$ (15)

### v2

@@ III-A Fault Model
Transistor miniaturization leads to more efficient computing systems, whereas its reliability side effects are becoming increasingly profound. With transistor scaling, soft error rates increase significantly, posing reliability concerns, in particular in the deployment of safety-critical applications [43, 44, 45]. Single Event Upset (SEU) due to soft errors may affect memory cells and flip a bit in a stored value. SEU in DNN accelerators affects either the parameters of a DNN (i.e., weights) or its activations (i.e., layers’ inputs), which are stored in memory elements such as registers or on-chip memories [19, 10].
In this paper, we build the mathematical analysis based on the single-bit error model in weights and activations separately. We assume that a single parameter or a single input to a layer is faulty, which is propagated to the output of the corresponding layer. By mathematical analysis, we model the fault propagation to the output of neurons and CNNs, resulting in calculating VF for them. In reliability analysis, it is not feasible to exploit the multiple-bit fault model since it requires a huge fault space to consider all the $2^{n}$ combinations, where $n$ is the number of bits and each bit can be in one of the two states: fault-free, or faulty. On the other hand, it is demonstrated that tests achieving full fault coverage on the single-bit fault model cover a vast majority of multiple-bit faults as well [46].
III-B Theoretical Single Fault Analysis in 32-bit Floating-Point
This subsection analyzes the effect of a single bitflip on the 32-bit floating-point data representation to comprehend how a value might change due to a single fault. The IEEE-754 32-bit floating-point data type is shown in Fig. 3. It contains 1 sign bit, 8 exponent bits, and 23 mantissa bits, and represents a number based on Eq. (2).
$\vskip-5.69054ptnumber=(-1)^{sign}\times 2^{E-127}\times(1+\sum_{i=0}^{i=23}{b_{23-i}\times 2^{i}})\vskip-8.53581pt$ (2)
@@ III-B Theoretical Single Fault Analysis in 32-bit Floating-Point
To model single bitflips, we consider the following lemmas:
  • Lemma 1: Any bitflip from 1 to 0 in a value decreases the value by $\epsilon$ while $\epsilon<0$, and any bitflip from 0 to 1 increases it by $\epsilon$ while $\epsilon>0$.
@@
      Sign: approximate $\epsilon$ as $\pm 2^{log_{2}(2\times x)}$.
@@
      Exponent: approximate $\epsilon$ as $\pm 2^{i-23}$, assuming that $x$ is a small value.
@@
      Mantissa: approximate $\epsilon$ as $\pm 2^{i}$.
    In all the cases, $\epsilon$ can be approximated as the nearest value of the power of two to the actual $\epsilon$. This approximation can lead to a unified representation for $\epsilon$. To that end, $\epsilon$ might be negative or positive (based on Lemma 1) and might be small or big (based on Lemma 2). Eventually, we can express it as a unified representation as shown in Eq. (4).
@@
  • Lemma 4: When a faulty value is used in a multiplication operation in CNNs, the erroneous output can also be approximated. It has been observed that the parameters in CNNs are mainly distributed around 0 and in the range of [1,1][-1,1] [49]. In order to approximate the error of multiplying two values when one of them is faulty, we can analyze it based on Eq. (5), where $x$ and $y$ are fault-free values and $x^{\prime}$ is the faulty value after a bitflip in $x$.
@@
III-C Single Bitflip Error Propagation in CNNs
We analyze the single bitflip error propagation in CNNs, considering its effect on the values of numbers. Each neuron in a convolutional (CONV) layer operates as shown in Eq. (6), where the Output Feature Map (OFMap) in $l$th layer and $k$th channel is obtained by the summation of multiplications between weights ($\hat{w}$) and Input Feature Map (IFMap, $\hat{x}$) plus bias ($b$). In CONV layers, $\hat{w}$ and $\hat{x}$ are 3-Dimensional (3D) $c_{in}\times n\times n$ arrays, where $c_{in}$ is the number of inputs channels to the layer, and $n$ is the kernel size.
@@ III-C Single Bitflip Error Propagation in CNNs
Figure 4: Fault propagation in the case of single bitflip in a filter.
@@
Based on the aforementioned theoretical analysis, to analyze the effect of single faults on CNNs, regardless of the fault occurrence in activations or weights, we need to identify an added value $\delta$ (as a power of two) at the output of the target neuron/channel where it changes the golden class of the CNN for each input image.
Figure 5: An overview of the DeepVigor+ approach for resilience analysis of CNNs.
@@ III-D DeepVigor+: Fault Resilience Analysis for DNNs
In this subsection, the steps of DeepVigor+ are presented. Fig. 5 illustrates the approach of the DeepVigor+ methodology for fault resilience analysis of CNNs against single faults. The method’s objective is to provide the Vulnerability Factor (VF) for layers and the entire CNN model when a single fault happens in activations or weights.
The inputs of the method are a pre-trained CNN and test data in a dataset. During the analysis, an OFMap or output channel is targeted, and multiple steps are taken to provide the VF metrics:
@@
    Vulnerability Values Identification: obtains vulnerability values for the target OFMap/channel,
@@
By obtaining the VF for the analyzed neurons and filters in a CNN, the VF for layers and the entire CNN can be derived. The details of each step are explained in the following. Note that each step presents a detailed description for analyzing a target neuron and then briefly links it to weights in filters.
Step 1 - Single Bitflip Analysis: This step conducts bitflips in the input activations of the target neuron to construct Vulnerability Values Search Space (VVSS) and Error Distribution Map (EDM) for it. VVSS is a set of approximated produced output errors, as in Eq. (9), by single bitflips in any bit locations of the target neuron’s inputs. EDM represents the distribution ratio of those approximated output errors.
Obtaining VVSS and EDM requires performing a single bitflip for each input and multiplying it by its corresponding weight as shown in Eq. (8). Nonetheless, it is an exhaustive operation with high time complexity. This complexity can be significantly reduced by using Algorithm 1. In this algorithm, for each bitflip, we obtain all possible exact $\delta$ values produced at the output of the neuron for all inputs at once.
Algorithm 1 One-Shot Bit-wise Error Analysis for a Neuron
@@
The faults in the input activations are independent for each input data. Hence, to produce all errors in a neural operation, we first flip the $i$th bit in all inputs (line 3) and then convert the binary representation to a value (line 4). Then, the difference of the erroneous_inputs with the fault-free input is computed (line 5), which represents all possible errors at inputs ($\epsilon$ in Eq. (8)) for all inputs added by a single bitflip in either of them. Then, a point-wise multiplication between the erroneous_inputs and weights results in producing all possible errors ($\delta$ in Eq. (9)) at the output (line 6).
Algorithm 1 produces all errors at the output of the target neuron at a single shot per bit. Thereafter, based on the presented theory in the subsections III-B and III-C, we generate the distribution of all produced errors as an Error Distribution Map (EDM) based on the Candidate Vulnerability Values (CVVs), which are a set of numbers as a power of two, shown in Eq. (11). Based on the experimental observations, we limit the CVV in the analysis between $2^{10}$ and $2^{10}$ for large values. The error values between $2^{-10}$ and $2^{-10}$ are discarded as their propagation effects on CNNs are negligible and masked.
@@
EDM contains the distribution ratio of the existing values over the values in output_errors_list in Algorithm 1, between each consecutive value in CVV, as shown in Eq. (12). EDM provides the ratio of error distribution for each candidate vulnerability value, demonstrating how much each CVV represents the produced errors at the output of the target neuron. Based on the EDM, each CVV whose distribution ratio is not zero is added to the Vulnerability Value Search Space (VVSS). This means that VVSS contains a subset of CVV that are the approximated error values produced by single faults in the inputs. Therefore, VVSS only includes the CVVs with a non-zero error distribution ratio.
@@
Step 2: Vulnerability Values Identification: This step exploits VVSS to identify the vulnerability values for the target neuron (i.e., OFMap) by exploring its constructed VVSS. The objective is to identify the maximum negative and minimum positive vulnerability values among the existing ones in VVSS for a neuron that changes the CNN from its golden classification for each input data.
@@
    $VVSS_{(-\infty,-1]}$: contains CVVs between $(-\infty,-1]$ that have a non-zero distribution ratio.
@@
    $VVSS_{(-1,0)}$: contains CVVs between (1,0)(-1,0) that have a non-zero distribution ratio.
@@
    $VVSS_{[1,+\infty)}$: contains CVVs between $1,+\infty)$ that have a non-zero distribution ratio.
@@
In box 3, a loss function $\mathcal{L}$ is calculated to obtain the effect of an added $\delta$ to the target OFMap on the output logits based on the summation of the differences between all output classes’ logits from the golden class. In this loss function, $\mathcal{E}_{c_{t}}$ represents the erroneous output logit of the golden top class, and $\mathcal{E}_{c_{j}}$ is the erroneous output logit of any other class.
In box 4, the gradient of loss function $\mathcal{L}$ w.r.t. the target OFMap is calculated to check if the CNN golden classification changes or not when the target OFMap is erroneous. If the gradient is 0, faults producing positive deviation in the corresponding neuron do not lead to a change of golden classification. In this case (box 5), the neuron’s vulnerability value is considered to be the biggest CVV assumed value for positive numbers (i.e., $2^{10}$) and the algorithm ends. If the gradient is not equal to 0, a value can be found for $\delta$ in the target OFMap that changes the golden classification of the CNN.
The rest of the algorithm attempts to find the minimum positive $\delta$ in VVSS. In box 6, the CNN’s classification results are examined when $\delta=1$ for the target neuron. It initializes the VVSS subset to be explored in the next steps. If the golden classification changes due to $\delta=1$, the vulnerability value is less than $1$, and we should explore $VVSS_{(0,1)}$ (box 7). Otherwise, $VVSS_{[1,+\infty)}$ should be explored (box 8). Then, $\delta$ is set to the middle element of the corresponding VVSS subset, and if it does not change the golden classification, the next bigger CVV should be explored. This process continues until the first value that changes the golden classification is found (loop 9). The final positive vulnerability value ($\delta^{+}$) for the target neuron is the minimum value in positive VVSS that does not change the golden class for each input.
To identify the maximum negative vulnerability value ($\delta^{-}$) that does not change the golden classification, the same procedure is carried out by exploring negative values in VVSS. As a result, the Vulnerability Value Range (VVR) for the corresponding neuron is obtained and is expressed in Eq. (13). VVR represents all induced error values in the outputs of a neuron, leading to a change of classification. Note that since bitflips in 32-bit floating-point might lead to large values, we assumed that any value larger than $2^{10}$ and less than $2^{-10}$ is critical for CNNs; therefore, these values represent $\infty$ and -$\infty$ respectively.
@@
The vulnerability values’ identification step for a target filter is similar to the one for a neuron. It is conducted similarly to the illustrated flowchart in Fig. 6 to obtain VVR for the corresponding output channel, resulting in VVR for each target channel in a CNN.
Step 3: Obtaining Vulnerability Factors: As mentioned, an Error Distribution Map (EDM) is obtained for the target neuron in step 1, representing the distribution ratio for each vulnerability value. On the other hand, the Vulnerability Value Range (VVR) is obtained in step 2. This step aims to map VVR to EDM to provide the VF for the target neuron, expressing the probability of changing the golden classification in the case of a single bitflip in its inputs.
Fig. 7 depicts how VF can be obtained by mapping VVR to an aggregated EDM. The aggregated EDM represents the summation of the distribution ratio for all values less than a CVV. As shown in Fig. 7, all values between $(\delta^{-},\delta^{+})$ are non-vulnerable, meaning that if an error deviates the OFMap as much as any value in this range, it will not change the golden output. Otherwise, the error value is vulnerable, leading to a change in the golden classification.
VF is obtained by subtracting the distribution ratio of the minimum positive vulnerability value ($\delta{+}$) and the maximum negative vulnerability value ($\delta^{-}$) as shown in Eq. (14). This equation expresses the portion of all errors produced at the output that are critical for the CNN classification, which is equivalent to the probability that a fault changes the CNN’s golden classification.
@@
The VF for a filter is obtained similarly. The EDM and VVR for the target filter are obtained in the previous steps. In this step, they are exploited to provide the VF for the target filter. Deriving the VF for individual neurons and filters leads to obtaining the VF for CONV layers and the entire CNN model.
To obtain the total MVF of the entire model Eq. (15) is introduced, where $L$ is the total number of layers in a CNN, $N_{l}$ and $W_{l}$ are the total number of output activations and weights in layer $l$ respectively, $LVF_{neuron_{l}}$ represents the Layer Vulnerability Factor (LVF) based on the average VF of the neurons in layer $l$, and $LVF_{filter_{l}}$ is the average VF of the filters in layer $l$. This equation is a layer-wise weighted average of LVFs throughout the layers of a CNN that includes both activations and filters vulnerability analysis, leading to the MVF for the entire model. It represents the probability of changing the golden classification if a fault occurs either in activations or weights.
@@
$\frac{\sum_{l=1}^{L}(\frac{N_{l}}{N_{l}+W_{l}}\times LVF_{neuron_{l}}+\frac{W_{l}}{N_{l}+W_{l}}\times LVF_{filter_{l}})}{L}$ (15)